# Targeted synthetic data generation for tabular data via hardness characterization

### v1

@@ 3.1 KNN Shapleys for hardness characterization: a toy example
(b) $1$NN Shapleys on the training data for $(x_{\mathrm{test}},y_{\mathrm{test}})\in\mathbb{R}\times\{0,1\},\ x_{\mathrm{train}}=0$.
(c) Expected value $\mathbb{E}(s_{\mathrm{train}})$ of the $1$NN Shapley for the data point $(x_{\mathrm{train}},0)$, for different values of $x_{\mathrm{train}}$.
Figure 2: Results for the toy experiment in Section 3.1, based on a 50-50 mixture of two univariate normals with unit variance and means $-1$ and $1$, and training data $\{(-1,0),(1,1),(x_{\mathrm{train}},0)\}$.
@@ 4 Results
In this section, the proposed hardness characterizer based on KNN Shapleys is first compared to the most common methods in the literature via comprehensive benchmarks on tabular data. Then, KNN Shapleys are calculated on a large credit default prediction dataset to identify the hardest points, which are later used to train and compare synthetic data generators. Lastly, the same pipeline is applied to a simulated dataset to verify consistency in the results.
@@ 4.1 Benchmark study
In Seedat et al. (2024), the problem of hardness characterization is approached quantitatively by comparing existing methods on how confidently they can identify different kinds of hard datapoints on a variety of OpenML datasets (Vanschoren et al., 2013). Their toolkit supports two tabular datasets (diabetes and cover) and it allows to perturb a chosen proportion $p$ of datapoints according to one of three hardness types (mislabeling, out-of-distribution, or atypical; cf. Section 2.1). For this study, we use $p\in\{0.05,0.1,0.15,0.2\}$ and the performance metric is the Area Under the Precision Recall Curve (AUPRC) of an MLP. Results are averaged across three independent runs per hardness type and then across the different hardness types. The KNN Shapleys-based characterizer is implemented by adding a custom class to the original GitHub repository (https://github.com/seedatnabeel/H-CAT). The results are plotted in Figure 3.
On the diabetes dataset, our characterizer outperforms existing methods for all choices of $p$, although based on the critical difference diagram in Figure 3(a), it can be noticed that performance is not significantly different from Data-IQ, the other method to offer native XGBoost support. On the cover dataset (see Figure 3(b)), our characterizer ranks 7th overall, although once again performance is statistically compatible with Data-IQ. Based on the mathematical intuition discussed in Section 3.1, KNN Shapleys are expected to be lower when a datapoint lies in a region of feature space where its target is atypical (cf. Section 2.1). When $p$ is too large relative to the dataset’s difficulty, this atypicality is compromised, causing a deterioration in the performance of our characterizer.
(a) Diabetes dataset
(b) Cover dataset
Figure 3: AUPRC heatmap and critical difference diagram for 13 different hardness characterizers (including KNN Shapleys, labelled knn) on the diabetes and cover datasets (Seedat et al., 2024).
Based on these results, low KNN Shapleys prove to be solid indentifiers of hardness. Most hardness characterization techniques in the literature were devised for neural networks, and do not natively support XGBoost. On the other hand, KNN Shapleys have the clear advantage over existing methods of not requiring model training and thus being deterministic and model agnostic. However, they require an external test set and can become expensive to calculate on very large datasets ($n>10^{6}$).
4.2 American Express dataset
@@
The Amex dataset has been first released as part of a Kaggle [^1] See https://www.kaggle.com/competitions/amex-default-prediction. competition, with winning submissions typically consisting in ensembles of different models. XGBoost was found to have the best performance on its own and is thus chosen as baseline for this study. In order to benefit from GPU acceleration, training was carried out on a cluster equipped with 4 GeForce RTX 3090 Ti GPUs and 256GB of memory. The number of gradient boosting rounds was set to an arbitrarily high number, using early stopping on validation Gini with a patience of 100 iterations of gradient boosting. Because of subsampling the datapoints and the features, the training process is non-deterministic and thus retraining a large number of times with different random seeds is required to mitigate randomness. This forces to choose a set of hyperparameters that offers low training times as well as solid validation performance. Through experimenting it was found, that setting the maximum tree depth to 4, the learning rate to 0.05, the subsampling ratio of datapoints to 0.8 and of features to 0.6, gives a validation Gini of $0.91986$ in under 2 minutes of training time.
4.2.1 Hardness characterization
@@
(a) Hardness characterization results
Data-IQ tag Easy Hard Ambiguous
% defaulters $19.07$ $58.73$ $51.09$
(d) Feature 0 distribution for defaulters and non-defaulters
Beyond these initial statistics, it is interesting to know where these points lie in feature space. By looking at the total reduction in loss due to a feature in the nodes where said feature is used to split the data, we can calculate the feature importances of a trained XGBoost classifier. In the case of the Amex dataset, the first feature dominates these scores, with an importance comparable to the bottom 100 features combined. Figure 4(d) displays the distribution of this feature separated by Data-IQ tag for both defaulters and non-defaulters: hard defaulters somewhat overlap with easy non-defaulters and vice-versa, while ambiguous defaulters are almost indistinguishable from ambiguous non-defaulters. This is a common issue with tabular data, referred to as outcome heterogeneity and is known to be captured by Data-IQ (Seedat et al., 2022).
KNN Shapleys are then computed for $K\in\{1,5,100\}$ under the assumption that the $n_{\mathrm{test}}$ available out-of-sample data points are representative of the entire data distribution. Figure 6 displays the distribution of 100NN Shapleys separated by Data-IQ’s tag: hard points as per Data-IQ exhibit low 100NN Shapleys, with very good adherence for the hardest points. In addition, the left tail of the ambiguous points lies in the region of low 100NN Shapleys and would thus be recognized as hard by our novel characterizer.
@@
Figure 6: Validation Gini after removing hardest points, with $95\%$ CIs.
To choose the best $K$, validation Gini is monitored as we re-train XGBoost after gradually removing the hardest datapoints. These points are expected to be the most valuable for XGBoost and thus the best hardness characterizer should be the one displaying the fastest drop in validation performance. 100NN Shapleys outperform both other choices of $K$ and Data-IQ (cf. Figure 6). This comparison highlights an issue with Data-IQ: after a sharp drop when the hard points are removed, performance decreases at a rate comparable to random as the most ambiguous points are removed. Seedat et al. (2022) argue that the removal of ambiguous points should actually make the model more robust, as aleatoric uncertainty is data-dependent and cannot be reduced unless additional features are collected. This is however not the case for the Amex dataset, and beyond the small amount of hard points identified, Data-IQ does not provide guidance on the next most valuable points.
4.2.2 Synthetic data generation
The $10\%$ hardest datapoints are chosen as training set for synthetic data generators, with the purpose of establishing whether augmenting only the hardest datapoints makes XGBoost more robust than non-targeted data augmentation. The choice of a hard threshold is imposed by the expensive computational cost of fine-tuning neural networks, and $10\%$ is justified as it is just before the elbow region where the decrease in validation performance when removing hard data slows down (cf. Figure 6). Both CTGAN and TVAE are implemented in Python using a heavily customized version of the sdv package (Patki et al., 2016), with extensive details about the training given in Appendix A.1. The quality of the synthetic samples can be examined directly for the first feature: Figure 7(a) (CTGAN) displays a good but not perfect overlap with real data, whereas from Figure 7(b) (TVAE) we can see better overlap of synthetic and real data with respect to CTGAN for the first feature.
(a) CTGAN
(b) TVAE
For models trained on the $10\%$ hardest points we augment by $100\%$, while for models trained using the entire dataset we augment by $10\%$, thus guaranteeing the same amount of synthetic samples across all experiments. Notice that for each attempt synthetic data has to be generated multiple times with different seeds to mitigate randomness and quantify uncertainty on the performance metric. The proposed routine, implemented in Python with the baytune package (Smith et al., 2019), requires very few attempts to find good proposals, and is thus more efficient than a grid search over a hyperparameter space, reducing the computational complexity of the problem.
(a) Scores in fine-tuning routine: hard points
(b) Scores in fine-tuning routine: entire dataset
(c) Variation in Gini after augmentation with $95\%$ CIs.
@@ 4.2.2 Synthetic data generation
(d) Scores for TVAE and TVAE on hard data points after $5\%,10\%,15\%,20\%$ augmentation, with $95\%$ confidence intervals estimated by generating synthetic data 10 times.
@@
Figures 8(a) and 8(b) display the scores after 20 attempts: the best scores are in both cases achieved by TVAE, and augmenting only the hardest points generally leads to a more significant improvement in validation performance, both for the best attempts and on average across different sets of hyperparameters. Information on the tuning grid and best hyperparameters are detailed in Appendix A.2.
@@
4.3 Simulation study
[…]
We proceed by tuning TVAE using the GPEI algorithm (described in Appendix A.1) both when augmenting the hardest $5\%$ by $100\%$ and when augmenting the entire dataset by $5\%$. Figures 9(c) and 9(d) display the scores after 10 attempts: the targeted approach results in a larger performance improvement, both for the best attempt and on average across different sets of hyperparameters. Information on the tuning process can be found in Appendix A.3. Finally, we augment by different amounts using the best attempts, with results displayed in Figure 9(b): hard points augmentation consistently outperforms the non-targeted approach and improves as more data is added.
(a) Scatter of training data (left) and hardest $5\%$ (right).
(c) Scores in fine-tuning routine: hard points
(d) Scores in fine-tuning routine: entire dataset
@@ 5 Conclusion and discussion
Methodologically, this work aims at bridging the gap between learning-based hardness characterization and game theoretic data valuation. The benchmarks reported in Section 4.1 represent the first quantitative comparison between a Shapley-based evaluator and existing hardness characterizers. While research on the two topics has often dealt with similar problems, such as mislabeling or data summarization (Paul et al., 2023; Jia et al., 2021), the two paths have never crossed.
Practically, this work attempts to change the perspective on data summarization: while existing literature has mostly focused on pruning the least valuable points from training data (Seedat et al., 2022; Kwon & Zou, 2023), here we propose augmenting the most valuable ones. We demonstrate that not all datapoints are created equal, and some play a more significant role in determining the predictive power of the final model: high-value data dictates which points we should either collect in larger amounts or, when this is not possible, synthetically generate.
Reproducibility statement
The Amex dataset is publicly available on the public data repository Kaggle. This work uses the denoised version by user raddar, which is publicly available on Kaggle in parquet format. The code to reproduce the results in this work is available in the GitHub repository tommaso-ferracci/Targeted_Augmentation_Amex. In particular, code to reproduce the results in Section 4 is available in the analyses/ folder. The outputs/ folder contains, in addition to figures and results, the model weights for the best synthesizers. Notice how every method from the ctgan and sdv packages uses the forked versions (tommaso-ferracci/CTGAN  and tommaso-ferracci/SDV) with the custom early stopping. Detailed instructions on how to create a Python virtual environment and install these dependencies are available in the Makefile. Details around hyperparameter tuning are also reported in Appendices A.1, A.2 and A.3.
Appendix A Appendix
@@ A.1 Training CTGAN and TVAE – Amex dataset
This section provides details about CTGAN and TVAE SDGs on the Amex dataset. A common issue when training GANs is the instability of both generator and discriminator losses. More specifically, the two typically move in opposite ways with oscillatory patterns, making it difficult to decide when to stop training. For this reason, we introduce a novel early stopping condition which tracks epoch after epoch a weighted average of the Kolgomorov-Smirnov (KS) statistic between real and synthetic data across all individual features. In particular, we choose to use as weights the feature importances, in order to focus more on the features relevant to XGBoost. The patience is set to $50$ epochs with a maximum number of epochs of $500$. If the early stopping condition is triggered, the model is reverted to the best epoch. A large batch size of 10 0001000010\,00010 000 is chosen to limit the number of updates per epoch and guarantee a smoother training process. Both the generator and discriminator are trained using the Adam optimizer (Kingma & Ba, 2017), setting the learning rate to $2\cdot 10^{-4}$, weight decay to $10^{-6}$, and the momentum parameters to $\beta_{1}=0.5$ and $\beta_{2}=0.9$. Figure 10(a) shows the losses and the statistic epoch by epoch: we can see that the losses move against each other and then more or less converge once the model cannot further improve. We notice oscillatory patterns in the tracked statistic, symptomatic of the training dynamics of the generator and discriminator pair, and the early stopping condition kicking in after around 100 epochs when the weighted KS statistic peaks at $0.83$.
Training of the VAE relies on the Adam optimizer with a learning rate of $10^{-3}$, weight decay of $10^{-5}$, $\beta_{1}$ of $0.9$ and $\beta_{2}$ of $0.999$. To avoid overfitting of training data and the subsequent generation of exact replicas of training points, we stop training once the maximum number of epochs is reached.
The following routine integrating both model selection and hyperparameter tuning is carried out separately for models trained on only the hardest points and models trained using the entire dataset:

### v2

@@ 3.1 KNN Shapleys for hardness characterization: an example
(a) $1$NN Shapleys on the training data for $(x_{\mathrm{test}},y_{\mathrm{test}})\in\mathbb{R}\times\{0,1\},\ x_{\mathrm{train}}=0$.
(b) Expected value $\mathbb{E}(s_{\mathrm{train}})$ of the $1$NN Shapley for the data point $(x_{\mathrm{train}},0)$, for different values of $x_{\mathrm{train}}$.
Figure 2: Results for the experiment in Section 3.1, based on a mixture of two normals with unit variance and means $-1$ and $1$, and training data $\{(-1,0),(1,1),(x_{\mathrm{train}},0)\}$.
@@ 4 Results
In this section, the performance of KNN Shapleys as hardness characterizer is first compared to the most common methods in the literature via benchmarks on tabular data. Second, we demonstrate the pipeline proposed in Section 3 on a simulated dataset. Finally, KNN Shapleys are calculated on a number of real-world tabular datasets to identify the hardest points, which are later used to train and compare synthetic data generators.
@@ 4.1 Benchmark study
Dataset
diabetes cover eye jannis
Hardness Proportion of perturbed points Proportion of perturbed points Proportion of perturbed points Proportion of perturbed points
characterizer 0.05 0.1 0.15 0.2 0.05 0.1 0.15 0.2 0.05 0.1 0.15 0.2 0.05 0.1 0.15 0.2
Agreement 0.951 0.927 0.902 0.877 0.954 0.932 0.908 0.885 0.951 0.927 0.903 0.878 0.951 0.927 0.903 0.879
AUM 0.956 0.934 0.912 0.890 0.965 0.947 0.928 0.908 0.953 0.932 0.909 0.885 0.961 0.942 0.922 0.902
CleanLab 0.956 0.933 0.911 0.888 0.967 0.950 0.930 0.911 0.954 0.933 0.911 0.887 0.961 0.942 0.922 0.902
Data-IQ 0.955 0.934 0.912 0.889 0.965 0.946 0.926 0.905 0.954 0.933 0.911 0.887 0.961 0.942 0.922 0.902
DataMaps 0.955 0.934 0.912 0.889 0.965 0.946 0.926 0.905 0.954 0.933 0.911 0.887 0.961 0.942 0.922 0.902
Detector 0.950 0.925 0.900 0.875 0.945 0.920 0.894 0.869 0.950 0.925 0.900 0.875 0.946 0.920 0.895 0.870
EL2N 0.942 0.914 0.886 0.859 0.922 0.890 0.861 0.832 0.945 0.915 0.886 0.860 0.926 0.894 0.863 0.833
Forgetting scores 0.951 0.925 0.900 0.874 0.951 0.927 0.902 0.877 0.951 0.926 0.901 0.875 0.951 0.927 0.902 0.877
GraNd 0.932 0.900 0.869 0.839 0.942 0.916 0.891 0.867 0.946 0.920 0.893 0.867 0.935 0.907 0.879 0.852
KNN-Shapleys 0.957 0.924 0.902 0.877 0.961 0.940 0.919 0.898 0.950 0.924 0.900 0.875 0.951 0.927 0.902 0.875
Prototypicality 0.953 0.927 0.904 0.880 0.968 0.950 0.929 0.909 0.953 0.932 0.909 0.885 0.959 0.939 0.918 0.897
Sample-loss 0.941 0.913 0.885 0.857 0.922 0.891 0.863 0.836 0.944 0.915 0.885 0.858 0.925 0.893 0.862 0.833
VoG 0.963 0.942 0.922 0.900 0.964 0.944 0.923 0.903 0.950 0.925 0.899 0.874 0.957 0.934 0.912 0.889
Table 1: AUPRC scores for 13 different hardness characterizers (including KNN Shapleys) on the diabetes, cover, eyes and jannis OpenML datasets, for different proportions of perturbed points.
@@
We demonstrate that KNN Shapleys have comparable performance to alternative hardness characterizers on tabular data. We provide a comparison with the results in the benchmarking study from Seedat et al. [2024], where the problem of hardness characterization is approached quantitatively by comparing existing methods on how confidently they can identify different kinds of hard datapoints on a variety of OpenML datasets [Vanschoren et al., 2013]. Their toolkit focuses on benchmarking four OpenML tabular datasets: diabetes, cover, eye and jannis. We compare KNN Shapleys to the following hardness characterization methods identified in Seedat et al. [2024]: Agreement [Carlini et al., 2019], AUM [Pleiss et al., 2020], CleanLab [Northcutt et al., 2021], Data-IQ [Seedat et al., 2022], DataMaps [Swayamdipta et al., 2020], Detector [Jia et al., 2023], EL2N [Paul et al., 2021], Forgetting scores [Toneva et al., 2019b], GraNd [Paul et al., 2021], Prototipicality [Sorscher et al., 2022], Sample-loss [Xia et al., 2022], and VoG [Agarwal et al., 2022]. In particular, for each dataset we perturb a chosen proportion $p$ of datapoints according to one of three hardness types (mislabeling, out-of-distribution, or atypical; cf. Section 2.1), with $p\in\{0.05,0.1,0.15,0.2\}$. We use the Area Under the Precision Recall Curve (AUPRC) of an MLP as the performance metric in order to facilitate direct comparisons, since most hardness characterizers analyzed in Seedat et al. [2024] are devised for neural networks and do not support XGBoost. The results are plotted in Table 1. Scores are averaged across three independent runs per hardness type and across hardness types.
Beyond the justifications described in Section 3, the results of the benchmarking study suggest that low KNN Shapleys are solid identifiers of hardness, with performances comparable with the most common techniques in literature. It must be emphasized that we do not claim that KNN Shapleys outperform other hardness characterizers, but we point out that the proposed scores achieve similar AUPRC to state-of-the-art methods, at a fraction of their computational cost. In particular, the main advantage lies in the fact that KNN Shapleys do not require training of any model, and they only require a calculation of a distance matrix. On the other hand, all the other methods in Table 1 require training of a model for calculation of the hardness characterization scores. This makes KNN Shapleys computationally efficient and model-agnostic, whereas most other methods require multiple model refitting procedures, rely on the training dynamics of neural networks, and it is not straighforward to extend them to gradient boosting for tabular data.
4.2 Simulation study
@@
We proceed by tuning TVAE using the GPEI algorithm (described in Appendix A.1) both when augmenting the hardest $\tau\cdot 100\%$ by $100\%$ and when augmenting the entire dataset by $\tau\cdot 100\%$, for $\tau\in\mathbb{R}_{+}$. Information on the tuning process can be found in Appendix A.3. Performance is measured via the normalized Gini coefficient $12\cdot\mathrm{AUCROC}-$ 1, with $\mathrm{AUCROC}$ denoting the Area Under the Receiver Operating Characteristic curve. Figure 3(b) displays the results for different values of $\tau$. The targeted approach results in a larger performance improvement, both for the best attempt and on average across different sets of hyperparameters. Hard points augmentation consistently outperforms the non-targeted approach and improves as more data is added.
(a) Scatter of training data (left) and hardest $5\%$ (right).
4.3 Case study: American Express data
@@
Hardness characterization
@@
KNN Shapleys are computed for $K\in\{1,5,100\}$ under the assumption that the $n_{\mathrm{test}}$ available out-of-sample data points are representative of the entire data distribution. Figure 4 displays the distribution of 100NN Shapleys separated by Data-IQ’s tag: hard points as per Data-IQ exhibit low 100NN Shapleys, with very good adherence for the hardest points. In addition, the left tail of the ambiguous points lies in the region of low 100NN Shapleys and would thus be recognized as hard by our novel characterizer. To choose the best $K$, validation Gini is monitored as we re-train XGBoost after gradually removing the hardest datapoints. These points are expected to be the most valuable for XGBoost and thus the best hardness characterizer should be the one displaying the fastest drop in validation performance. 100NN Shapleys outperform both other choices of $K$ and Data-IQ (cf. Figure 5). This comparison highlights an issue with Data-IQ: after a sharp drop when the hard points are removed, performance decreases at a rate comparable to random as the most ambiguous points are removed. Seedat et al. [2022] argue that the removal of ambiguous points should actually make the model more robust, as aleatoric uncertainty is data-dependent and cannot be reduced unless additional features are collected. This is however not the case for the Amex dataset, and beyond the small amount of hard points identified, Data-IQ does not provide guidance on the next most valuable points.
Figure 5: Validation Gini after removing the hardest points with KNN Shapleys, Data-IQ, and randomly, with $95\%$ CIs.
Synthetic data generation
The $10\%$ hardest datapoints are chosen as training set for synthetic data generators, with the purpose of establishing whether augmenting only the hardest datapoints makes XGBoost more robust than non-targeted data augmentation. The choice of a hard threshold is imposed by the expensive computational cost of fine-tuning neural networks, and $10\%$ is justified as it is just before the elbow region where the decrease in validation performance when removing hard data slows down (cf. Figure 5). Details about the training of the SDGs are given in Appendix A.1.
For models trained on the $10\%$ hardest points, we augment by $100\%$, while for models trained using the entire dataset we augment by $10\%$, thus guaranteeing the same amount of synthetic samples across all experiments. Notice that for each attempt synthetic data has to be generated multiple times with different seeds to mitigate randomness and quantify uncertainty on the performance metric.
(a) Variation in Gini after augmentation with $95\%$ CIs.
@@
(b) Scores for TVAE and TVAE on hard data points after $5\%,10\%,15\%,20\%$ augmentation, with $95\%$ confidence intervals estimated by generating synthetic data 10 times.
@@
The best scores are in both cases achieved by TVAE, and augmenting only the hardest points generally leads to a more significant improvement in validation performance, both for the best attempts and on average across different sets of hyperparameters. Information on the tuning grid and best hyperparameters are detailed in Appendix A.2.
@@
4.4 Additional results
Dataset Augmenting hardest 5% by 200% Augmenting hardest 10% by 100% Augmenting hardest 20% by 50% Augmenting by 10% (non targeted)
diabetes 0.00084 0.00050 0.00193 0.00159
credit_card 0.00429 0.00528 0.00173 0.00303
blobs 0.00049 -0.00130 -0.00127 -0.00255
support 0.00229 0.00379 0.00716 0.00592
Table 2: Variation in Gini after 10% data augmentation for targeted and non-targeted approaches on four different datasets.
To extend the results beyond the case study on the Amex dataset, and further verify the robustness of the proposed pipeline, the same hardness characterization and data augmentation pipeline are applied to other common binary classification datasets from the UC Irvine (UCI) machine learning repository. Similarly to the Amex dataset, hyperparameter tuning of CTGAN and TVAE is performed at the synthetic data generation stage, and the randomness of both XGBoost training and data augmentation is mitigated by taking the mean over multiple runs. In addition, the smaller size of these datasets allows to experiment with different thresholds for hardness characterization, specifically 5%, 10% and 20%. The results are reported in Table 2. The results demonstrated that targeted augmentation outperforms the non-targeted approach for all datasets, with the choice of the threshold $\tau$ proving critical.
@@ 5 Conclusion and discussion
Methodologically, this work aims at bridging the gap between learning-based hardness characterization and game theoretic data valuation. The benchmarks reported in Section 4.1 represent the first quantitative comparison between a Shapley-based evaluator and existing hardness characterizers. While research on the two topics has often dealt with similar problems, such as mislabeling or data summarization [Paul et al., 2023, Jia et al., 2021], the two paths have never crossed.
Practically, this work attempts to change the perspective on data summarization: while existing literature has mostly focused on pruning the least valuable points from training data [Seedat et al., 2022, Kwon and Zou, 2023], here we propose augmenting the most valuable ones. We demonstrate that not all datapoints are created equal, and some play a more significant role in determining the predictive power of the final model: high-value data dictates which points we should either collect in larger amounts or, when this is not possible, synthetically generate.
Details on reproducibility
The Amex dataset is publicly available on the public data repository Kaggle. This work uses the denoised version by user raddar, which is publicly available on Kaggle in parquet format. The code to reproduce the results in this work is available in the GitHub repository tommaso-ferracci/Targeted_Augmentation_Amex. In particular, code to reproduce the hardness characterization and data augmentation pipeline on both the Amex dataset and the benchmark datasets is available in the analyses/ folder. The outputs/ folder contains, in addition to figures and results, the model weights for the best synthesizers. Every method from the ctgan and sdv packages uses forked versions (tommaso-ferracci/CTGAN  and tommaso-ferracci/SDV) with custom early stopping. Detailed instructions on how to install these dependencies are available in the Makefile. Details around training are also reported in Appendices A.1, A.2 and A.3.
APPENDIX
Targeted synthetic data generation for tabular data via hardness characterization
Appendix A Training details
@@ A.1 Training CTGAN and TVAE – Amex dataset
In the following, we provide details about CTGAN and TVAE SDGs on the Amex dataset. A common issue when training GANs is the instability of both generator and discriminator losses. More specifically, the two typically move in opposite ways with oscillatory patterns, making it difficult to decide when to stop training. For this reason, we introduce a novel early stopping condition which tracks epoch after epoch a weighted average of the Kolgomorov-Smirnov (KS) statistic between real and synthetic data across all individual features. In particular, we choose to use as weights the feature importances, in order to focus more on the features relevant to XGBoost. The patience is set to $50$ epochs with a maximum number of epochs of $500$. If the early stopping condition is triggered, the model is reverted to the best epoch. A large batch size of 10 0001000010\,00010 000 is chosen to limit the number of updates per epoch and guarantee a smoother training process. Both the generator and discriminator are trained using the Adam optimizer [Kingma and Ba, 2017], setting the learning rate to $2\cdot 10^{-4}$, weight decay to $10^{-6}$, and the momentum parameters to $\beta_{1}=0.5$ and $\beta_{2}=0.9$. Figure 7(a) shows the losses and the statistic epoch by epoch: we can see that the losses move against each other and then more or less converge once the model cannot further improve. We notice oscillatory patterns in the tracked statistic, symptomatic of the training dynamics of the generator and discriminator pair, and the early stopping condition kicking in after around 100 epochs when the weighted KS statistic peaks at $0.83$.
Training of the VAE relies on the Adam optimizer with a learning rate of $10^{-3}$, weight decay of $10^{-5}$, $\beta_{1}$ of $0.9$ and $\beta_{2}$ of $0.999$. To avoid overfitting of training data and the subsequent generation of exact replicas of training points, we stop training once the maximum number of epochs is reached. The following routine integrating both model selection and hyperparameter tuning is carried out separately for models trained on only the hardest points and models trained using the entire dataset: